OpenStudy (anonymous):

Given 5x^3 - 17x^2 + 12x -6 = (x^2-3x+1) (5x +A) + Bx + C for all values of x, a) find the values of A B C ans: A= -2 B = 1 C = -4 b) HENCE, find the remainder when 5x^3 -17 x ^2 + 12x -6 is divided by (x^2-3x + 1 )

mathslover (mathslover):

k wait ..

mathslover (mathslover):

So may i know the exact question please

mathslover (mathslover):

@Sgstudent please give us the exact question

OpenStudy (anonymous):

This is the exact question ? lol

mathslover (mathslover):

I think u have typed wht u did instead of the exact question ..

OpenStudy (anonymous):

okay for all values of x is after the +C

OpenStudy (anonymous):

um.. u already got c=-4 .: the remainder is x-4. recognise that the original equation is Dividend=divisor*quotient + remainder where remainder is Bx+C for all values of x just means that the equality is true for all values..

OpenStudy (anonymous):

but i need to show working my dear

OpenStudy (anonymous):

compare co efficients form eqn's solve them

OpenStudy (anonymous):

i dont get the hence part

OpenStudy (anonymous):

what.. just divide the dividend by the divisor.. you will get 5x-2 + x-4/x^2-3x+1 where 5x-2 = 5x+A and x-4=Bx+C..

OpenStudy (anonymous):

i dont think i learn that yet any other way?

OpenStudy (anonymous):

read http://en.wikipedia.org/wiki/Polynomial_long_division DIVISION TRANSFORMATION PART .. youre looking for the R(x)

OpenStudy (anonymous):

what do you get ?

OpenStudy (anonymous):

.. as stated in the question R(x) = Bx + C so the remainder is that. what year level are u in?

OpenStudy (anonymous):

@Sgstudent simplify the right hand side of equation (x^2-3x+1) (5x +A) + Bx + C get the standard form of it then compare the corresponding coefficients of 2 polynomial